OpenStudy (anonymous):

10 + squareroot 10m-1=13

OpenStudy (anonymous):

any choices

OpenStudy (anonymous):

3 = sqrt(10m - 1) 9 = 10m -1 10 = 10m 1 = m

OpenStudy (superdavesuper):

I think the question needs to more clear: is it 10+SQRT(10m-1)=13 or 10 + SQRT(10m) -1=13?

OpenStudy (anonymous):

it's obviously sqrt(10m-1) not sqrt(10m) -1 since the latter would result in a very strange... uhmm... result >.>

OpenStudy (superdavesuper):

@snuggles7 haha true but I have seen stranger questions n answers before! Back to the question, first thing is to subtract 10 on both side of =. then square to get rid of SQRT...